OpenStudy (anonymous):

find the length of the side of a square with an area of 144 48 12 206 288

OpenStudy (anonymous):

Hi Welcome to Openstudy!! First we have to know the formula for the area of a square. do you know the formula?

OpenStudy (anonymous):

um let me check

OpenStudy (anonymous):

A= LW or Sr

OpenStudy (anonymous):

so the formula for area of a square is \(\sf A=s^2\) where s is the side length and A is the formula. Now you are given the Area and not the side, so you have to solve for the side length. \(\sf A=s^2\\144=s^2\) in order to solve for s, we have to get rid of the exponent 2, in order to get rid of the exponent 2, we have to take the square root of both sides... can you do that for me? :)

OpenStudy (anonymous):

\(\sf A=s^2\) is the formula because the length and the width of a square are equal, right? ^_^

OpenStudy (anonymous):

or LW

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yes, area is Length times Width... but in a square, length and width are equal.. because all sides of a square are equal to each other

OpenStudy (anonymous):

oh yay your right

OpenStudy (anonymous):

ok but whats the length and width 6 if yay then times 6x6 = 36

OpenStudy (anonymous):

its not any of the answers though

OpenStudy (anonymous):

where did you get 6 from? have you done what I have discussed above?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you just plz tell me the answer

OpenStudy (anonymous):

Sorry, we're not supposed to give answers, we're just here to guide you. I'm just following the rules. \(\sf A=s^2\\144=s^2\\s=\sqrt{144}\\s=?\) now, evaluate that ^ . What value you'll get for the side length (s)?

OpenStudy (anonymous):

@AWESOME...A do you have a calculator with you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have the square root function?

OpenStudy (anonymous):

is it B

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

now you get it :) @AWESOME...A

OpenStudy (anonymous):

cool